Gabby07:

A rectangular garden measures 32ft by 44ft. Surrounding (and bordering) the garden is a path 4ft wide. Find the area of this path. Be sure to include the correct unit in your answer.

thoughfulsky:

Ok

thoughfulsky:

We're going to do this step by step

Gabby07:

okkk

thoughfulsky:

This is the rectangle without garden

thoughfulsky:

so area of a rectangle = l*b

thoughfulsky:

Therefore, area of rectangle = 44*32 = 1408 ft

Gabby07:

ok but how do i find the 4 ft path

thoughfulsky:

Wait

thoughfulsky:

So the question wants the area of the garden which can be done by subtracting: area of the rectangle along with the garden- area of just the rectangle

thoughfulsky:

so here's the rectangle with the garden. So whats the measures of the rectangle with the garden?

thoughfulsky:

therefore u see with the garden the measure of the rectangle increases by 4ft on all the four sides

thoughfulsky:

Length of the rectangle along with the garden = 32+4 = 36 ft

thoughfulsky:

Breadth of the rectangle along with the garden = 44+4 =48 ft

thoughfulsky:

Area of the rectangle along with the garden = l*b =36*48 =1728 ft

thoughfulsky:

Area of the garden = area of the rectangle along with Gardn- area of rectangle = 1728-1420 =308 ft

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky Length of the rectangle along with the garden = 32+4 = 36 ft \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky Breadth of the rectangle along with the garden = 44+4 =48 ft \(\color{#0cbb34}{\text{End of Quote}}\) Take a look at your drawing once more, it should be +8

thoughfulsky:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky Length of the rectangle along with the garden = 32+4 = 36 ft \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky Breadth of the rectangle along with the garden = 44+4 =48 ft \(\color{#0cbb34}{\text{End of Quote}}\) Take a look at your drawing once more, it should be +8 \(\color{#0cbb34}{\text{End of Quote}}\) What?

AZ:

The new dimensions of the rectangle when you're adding in the path

thoughfulsky:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ The new dimensions of the rectangle when you're adding in the path \(\color{#0cbb34}{\text{End of Quote}}\) What??

AZ:

The dimensions you found including the path is incorrect. There is a 4 foot path on both sides so it should be 32 + 4 + 4 = 40 and 44 + 4 + 4 = 52

1 attachment
thoughfulsky:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ The dimensions you found including the path is incorrect. There is a 4 foot path on both sides so it should be 32 + 4 + 4 = 40 and 44 + 4 + 4 = 52 \(\color{#0cbb34}{\text{End of Quote}}\) Are u sure 🤨

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky Are u sure 🤨 \(\color{#0cbb34}{\text{End of Quote}}\) Just look at your own drawing???

thoughfulsky:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky Are u sure 🤨 \(\color{#0cbb34}{\text{End of Quote}}\) Just look at your own drawing??? \(\color{#0cbb34}{\text{End of Quote}}\) My drawing shows the equal width all around. But for area u only take the length of one side and breadth one side so +4 not +8.

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky My drawing shows the equal width all around. But for area u only take the length of one side and breadth one side so +4 not +8. \(\color{#0cbb34}{\text{End of Quote}}\) It's a 4 foot path on EACH side (just like how you drew it) and I am well aware how you calculate the area the new length dimension would be adding 4 feet since there's a 4 foot path on each side as well as the width would be +4 + 4 since there's an extension by 4 feet on each side

thoughfulsky:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @thoughfulsky My drawing shows the equal width all around. But for area u only take the length of one side and breadth one side so +4 not +8. \(\color{#0cbb34}{\text{End of Quote}}\) It's a 4 foot path on EACH side (just like how you drew it) and I am well aware how you calculate the area the new length dimension would be adding 4 feet since there's a 4 foot path on each side as well as the width would be +4 + 4 since there's an extension by 4 feet on each side \(\color{#0cbb34}{\text{End of Quote}}\) Bye. I also know how to calculate ig e.e